OpenStudy (anonymous):

Nadia took two practice exams to prepare for her college entrance exam. On the first practice exam, she scored 1,100. After studying, her score increased by 25%. What was her score on the second practice exam? 1,125 1,350 1,375 1,467

OpenStudy (anonymous):

@timo86m HELP PLS

OpenStudy (anonymous):

Is one thousand one hundred =1100 ?

OpenStudy (anonymous):

x+(.25*x)= score on second practice Where x is the old score.

OpenStudy (anonymous):

@timo86m yes 1,100 and ok

OpenStudy (anonymous):

Plug in 1100 in every x then :D 1,100+(.25* 1,100) can you evaluate?

OpenStudy (anonymous):

Ill try..

OpenStudy (anonymous):

Hint do what is inside the parenthesis first. Due to order of operations :D

OpenStudy (anonymous):

\[1,100 + (25 \times 1,100) \] \[1,100 + 275 \] \[= 1375\]

OpenStudy (anonymous):

@timo86m thannk you for helping me (:

OpenStudy (anonymous):

1375 is right and NP Julietta :D